OpenStudy (anonymous):

Given f(x) = 5x - 6 and g(x) = x2 - 5x + 6, find the function (f - g)(x). Use the ^ key for the exponent. Enter your answer as the following example: -2x^2+3x+10.

OpenStudy (ghdoru11):

x2 or x^2?

OpenStudy (ghdoru11):

this is not the answer

OpenStudy (ghdoru11):

g(x)=x2... or x^2...?

OpenStudy (anonymous):

idk im really confussed where to even start

OpenStudy (mathmale):

f(x) = 5x - 6 and g(x) = x2 - 5x + 6 are separate functions. Your job is to subtract the 2nd function from the first; we know this because we're supposed to evaluate (f-g)(x). Write f(x) = 5x - 6 -g(x) = -( x2 - 5x + 6) ---- ------------ Here we have four columns. The first column can be written as f(x)-g(x). Find the sums of the second, third and fourth columns.

OpenStudy (mathmale):

Note that f(x) = 5x - 6 -g(x) = -( x2 - 5x + 6) ---- ------------ simplifies to f(x) = 5x - 6 -g(x) = - x^2 + 5x - 6 ---- ------------- Please finish the combining.